OpenStudy (anonymous):

A retailer has two options for series of trade discounts: 1: If retailer can purchase 30 total items, the retailer receives a 20% discount on the entire purchase. If retailer can purchase even more (say 50 items), retailer will receive an additional 30% discount on the previously discounted price. 2: If retailer can purchase 30 total items, retailer will receive a 30% discount on the entire purchase. If retailer can purchase even more (say 50 items), retailer will receive an additional 20% discount on the previously discounted price. Which is better deal?

OpenStudy (anonymous):

Wouldn't they both be the same since no net price is given?

OpenStudy (texaschic101):

1. lets just say the items are $2 each....purchasing 30 of them would be $60 with a discount of 20% which is (.20)(60) = 12.....60 - 12 = $48 would be the cost That means he can get 30 items for $48. 50 items at $2 each = $100. And an additional 30% discount added to the 20% discount = 50% discount. 50% discount off of 100 = $50 That means he can purchase 50 items for $50. So would you rather get 30 items for $48..or spend $2 more and get 20 more items ? 2. 30 items at $2 each with 20% discount = $48. 50 items at $2 each = $100. 20% discount + another 20% = 40% off. (.40)(100) = 40. $100 - $40 = $60. So he purchased 50 items for $60. So we have 30 items for $48.... 30/48 = 50/x (30 items for $48 = 50 items for x bucks) 30x = (48)(50) 30x = 2400 x = 80 at this rate, purchasing 50 items would cost $80 or... 50 items for $60. which one would you choose ?